OpenStudy (anonymous):

Evaluate –x2 + 2x – 4 with x = –2. i keep getting 4 as my ans but thats not an option!

OpenStudy (anonymous):

positive 2

OpenStudy (anonymous):

(-2)^2 +2(-2) -4 4 +-4 - 4 -4

OpenStudy (anonymous):

can be solved multiply ways, depends on how you treat the leading negative sign...what are your options

OpenStudy (anonymous):

8 -8 12 -12 those are my options

OpenStudy (anonymous):

-x2+2x-4 x= -2 -(-2)*2+2(-2)-4 = 8

OpenStudy (smokeysthename):

got -12

OpenStudy (smokeysthename):

i got -12

OpenStudy (anonymous):

-12 is right

OpenStudy (smokeysthename):

-(-2)^2+2(-2)-4 -2^2 =4 -4+2(-2)-4 2(-2)=-4 -4-4-4=-12

OpenStudy (anonymous):

\[-(-2)^2+2(-2)-4 \]\[-4+-4-4 \]\[-4-4-4 \]\[-12 \]

OpenStudy (anonymous):

i got -12 the first time LOL